OpenStudy (anonymous):

what is the product of (5 sqrt 5)(6 sqrt 4) simplify your answer a.30 sqrt9 b.60 sqrt 4 c. 60 sqrt 24 d. 150 sqrt 4 I believe it is A

OpenStudy (nysha):

i don't know I'm sorry i wish i could help

OpenStudy (anonymous):

ok then I strongly agree with B

OpenStudy (anonymous):

I meant C

OpenStudy (anonymous):

anyone know?

OpenStudy (rushwr):

\[5\sqrt{5} * 6\sqrt{4}\] is this ur question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

rushwr

OpenStudy (anonymous):

@nightcore144

OpenStudy (anonymous):

@Owl Buzz

OpenStudy (rushwr):

\[5\sqrt{5} * 6\sqrt{4} = \sqrt{25*5 * 36 *4} = \sqrt{18000}\]

OpenStudy (anonymous):

Rushwr?

OpenStudy (rushwr):

\[\sqrt{18000} = \sqrt{9 * 2 * 100*10}\]

OpenStudy (anonymous):

I don't know A, B, C, or is that D?

OpenStudy (rushwr):

I dnt get any answers given ! I got 60 sqr 5

OpenStudy (anonymous):

im sorry i cant help you

OpenStudy (anonymous):

ok nightcore

OpenStudy (anonymous):

so I should go with B rushwr

OpenStudy (anonymous):

it was between the first three choices for me

OpenStudy (anonymous):

so I was thinking C I thought it was simple math but then I saw B too

OpenStudy (rushwr):

not soo sure because i didnt get the answer correct. Are u sure about the question u gave me ? because u dnt write 6sqr4 like this because square root of 4 is 2 so u write it as 12 instead.

OpenStudy (anonymous):

( 5 squared 5) (6 squared 4)

OpenStudy (anonymous):

I'm sure that's the question your also suppose to simplify it

OpenStudy (rushwr):

This is completely different from what u said earlier u said square root earlier.

OpenStudy (rushwr):

\[5^{2} * 5 * 6^{2} * 4\]

OpenStudy (anonymous):

yeah it is using sqrt

OpenStudy (anonymous):

not that new equation you just posted

OpenStudy (anonymous):

so I'l be going with B I guess